# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.1 Proofs of the Auxiliary Lemmas and Resutls
We first compute a bound for VAR(1) case. Then use it for VAR(p). From Lütkepohl [2005], we get $\bm{\Gamma}(j)=\sum_{i=0}^{\infty}\bm{A}^{j+i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}=\bm{A}^{j}\sum_{i=0}^{\infty}\bm{A}^{i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}$ for a VAR(1) process, where $\bm{A}=\bm{\Gamma}(1)\bm{\Omega}$.
Using the telescopic sum setting $\bm{\Gamma}(0)-\bm{A}\bm{\Gamma}(0)\bm{A}=\bm{C}_{1}$, we have $\sum_{i=0}^{\infty}\bm{A}^{i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}=\bm{\Gamma}(0)-\lim_{k\rightarrow\infty}\bm{A}^{k}\bm{\Gamma}(0)(\bm{A}^{k})^{\mathrm{T}}=\bm{\Gamma}(0)$ under our parametrization. Hence, $\bm{\Gamma}(j)=\bm{A}^{j}\bm{\Gamma}(0)$.
$\bm{\Gamma}(1)=\bm{U}_{1}\bm{V}_{1}^{\mathrm{T}}$ and $\bm{C}_{1}=\bm{\Omega}^{-1}-\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}}$. We have $\bm{V}_{1}\bm{\Omega}\bm{V}_{1}^{\mathrm{T}}=\bm{I}$ and $\bm{\Omega}^{-1}\geq\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}}$. Alternatively, $\lambda_{\min}(\bm{\Omega}^{-1}-\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}})\geq 0$.
Since, $\bm{V}_{1}^{\mathrm{T}}\bm{\Omega}\bm{V}_{1}=\bm{I}$, we have $\|\bm{V}_{1}^{\mathrm{T}}\bm{\Omega}^{1/2}\|^{2}_{\mathrm{op}}=1$. Hence $\|\bm{V}_{1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Omega}^{-1/2}\|^{2}_{\mathrm{op}}$
$\bm{A}^{k}-\bm{B}^{k}=\bm{A}^{k-1}(\bm{A}-\bm{B})+(\bm{A}^{k-1}-\bm{B}^{k-1})\bm{B}$. Recursively using this relation and applying norm, we have $\|\bm{A}^{k}-\bm{B}^{k}\|^{2}_{F}\leq\|\bm{A}-\bm{B}\|^{2}_{F}(\sum_{i=1}^{k}\|\bm{A}\|_{\mathrm{op}}^{2(k-i)}\|\bm{B}\|_{\mathrm{op}}^{2(i-1)})$, using the triangle inequality and the submultiplicative property of operator norm.
After some simplification using GP-series sum results, $\sum_{i=1}^{k}\|\bm{A}\|_{\mathrm{op}}^{2(k-i)}\|\bm{B}\|_{\mathrm{op}}^{2(i-1)}=\frac{\|\bm{A}\|_{\mathrm{op}}^{2k}-\|\bm{B}\|_{\mathrm{op}}^{2k}}{\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}}=\frac{\{\|\bm{B}\|^{2}_{\mathrm{op}}+(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})\}^{k}-\|\bm{B}\|_{\mathrm{op}}^{2k}}{\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}}=\sum_{i=0}^{k-1}{k\choose i}\|\bm{B}\|^{i}_{\mathrm{op}}(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})^{k-1-i}$.
Hence, $\|\bm{A}^{k}-\bm{B}^{k}\|^{2}_{F}\leq\|\bm{A}-\bm{B}\|^{2}_{F}\left\{\sum_{i=0}^{k-1}{k\choose i}\|\bm{B}\|^{i}_{\mathrm{op}}(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})^{k-1-i}\right\}\leq k\|\bm{A}-\bm{B}\|^{2}_{F}\min\{\|\bm{A}\|_{\mathrm{op}}^{2(k-1)},|\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}|^{k-1}\}$
$\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}}+\sum_{k=1}^{T-1}(T-k)\|\bm{\Upsilon}_{1}(k)-\bm{\Upsilon}_{2}(k)\|^{2}_{F}$.
$\|\bm{\Upsilon}_{1}(k)-\bm{\Upsilon}_{2}(k)\|^{2}_{F}\leq\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}^{k}-\bm{A}_{2}^{k}\|_{\mathrm{F}}+\|\bm{A}_{2}\|^{k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}$.
$\|\bm{A}_{1}-\bm{A}_{2}\|_{\mathrm{F}}\leq\|\bm{\Upsilon}_{1}(1)\|^{2}_{\mathrm{op}}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{F}+\|\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(1)-\bm{\Upsilon}_{2}(1)\|^{2}_{F}\leq\|\bm{U}_{1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1}\|^{2}_{\mathrm{op}}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{F}+\|\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}\{\|\bm{U}_{1,1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,1}-\bm{V}_{1,2}\|^{2}_{F}+\|\bm{V}_{1,2}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,1}-\bm{U}_{1,2}\|^{2}_{F}\}$.
And,
$\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$
$\quad\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}}+\sum_{k=1}^{T-1}(T-k)\{\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}^{k}-\bm{A}_{2}^{k}\|^{2}_{\mathrm{F}}+\|\bm{A}_{2}\|^{k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\}$
$\quad\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}}+\sum_{k=1}^{T-1}(T-k)\{\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}k\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{F}a^{k-1}$
$\qquad+\|\bm{A}_{2}\|^{2k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\}$
$\quad\leq(T\frac{b-b^{\mathrm{T}}}{1-b}-\frac{b\{(T-1)b^{\mathrm{T}}-Tb^{T-1}+1\}}{(1-b)^{2}}+T)\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}$ (A.5)
$\qquad+T\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{F}\sum_{k=1}^{T-1}ka^{k-1}$
$\quad\leq(\frac{Tb-Tb^{2}+b^{T+1}-b}{(1-b)^{2}}+T)\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}$
$\qquad+T\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{F}\frac{1+(T-2)a^{T-1}-(T-1)a^{T-2}}{(1-a)^{2}}$ (A.6)
where $a=\|\bm{A}_{1}\|^{2}_{\mathrm{op}}$ and $b=\|\bm{A}_{2}\|^{2}_{\mathrm{op}}$. Since, $a<1$, we have $\frac{1+(T-2)a^{T-1}-(T-1)a^{T-2}}{(1-a)^{2}}<\frac{1}{(1-a)^{2}}$. This comes in terms of the model parameters.
@@
It is easy to verify that $\frac{1+(T-2)a^{T-1}-(T-1)a^{T-2}}{(1-a)^{2}}$ and $\frac{Tb-Tb^{2}+b^{T+1}-b}{(1-b)^{2}}$ are increasing in $a$ and, $b$ respectively. Following are the limits,
@@
Due to VAR(1) representation of VAR$(p)$ [Ghosh et al., 2019], Equation (A.6) and $\bm{A}_{\ell}=\bm{\kappa}_{\ell,p}\bm{\Upsilon}_{\ell,p-1}^{-1}$ for $\ell=1,2$, we have
@@
$\leq T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}$
$\qquad+T^{3}\|\bm{\Upsilon}_{1,p-1}\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{F}$
$\leq T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}$
$\qquad+T^{3}\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{F}$
Proof of Result 11.
$\bm{D}_{j}=\bm{D}_{j-1}-\bm{W}_{j}^{\mathrm{T}}\bm{C}_{j-1}^{-1}\bm{W}_{j}$ with $\bm{W}_{j}=\bm{U}_{j}\bm{V}_{j}$.
$\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{F}+\|\bm{W}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{W}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{F}$.
We have $\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{F}\leq\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}}$ and $\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}$, $\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}$.
We will be able to use mathematical induction to show that the difference can be bounded by the parameters.
Thus, $\|\bm{\Gamma}^{-1}_{1,j}-\bm{\Gamma}^{-1}_{2,j}\|^{2}_{F}\leq\|\bm{G}_{1}-\bm{G}_{2}\|_{\mathrm{F}}+(\|\bm{G}_{1}\|^{2}_{\mathrm{op}}+\|\bm{G}_{2}\|^{2}_{\mathrm{op}})\|\bm{H}_{1}-\bm{H}_{2}\|^{2}_{F}\leq\|\bm{\Gamma}^{-1}_{1,j-1}-\bm{\Gamma}^{-1}_{2,j-1}\|^{2}_{F}+\|\bm{D}^{-1}_{1,j}-\bm{D}^{-1}_{2,j}\|^{2}_{F}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}$.
$\|\bm{D}^{-1}_{1,j}-\bm{D}^{-1}_{2,j}\|^{2}_{F}\leq\|\bm{D}^{-1}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{D}^{-1}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}\leq\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}$.
since $\|\bm{\Upsilon}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{1,p}^{-1}\|^{2}_{\mathrm{op}},\|\bm{\Upsilon}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}$.
${\displaystyle\mathbf{P}={\begin{bmatrix}\mathbf{A_{1}}&\mathbf{A_{2}}\\ \mathbf{A_{3}}&\mathbf{A_{4}}\end{bmatrix}}^{-1}={\begin{bmatrix}\left(\mathbf{A_{1}}-\mathbf{A_{2}A_{4}}^{-1}\mathbf{A_{3}}\right)^{-1}&-\left(\mathbf{A_{1}}-\mathbf{A_{2}A_{4}}^{-1}\mathbf{A_{3}}\right)^{-1}\mathbf{A_{2}A_{4}}^{-1}\\ -\mathbf{A_{4}}^{-1}\mathbf{A_{3}}\left(\mathbf{A_{1}}-\mathbf{A_{2}A_{4}}^{-1}\mathbf{A_{3}}\right)^{-1}&\quad\mathbf{A_{4}}^{-1}+\mathbf{A_{4}}^{-1}\mathbf{A_{3}}\left(\mathbf{A_{1}}-\mathbf{A_{2}A_{4}}^{-1}\mathbf{A_{3}}\right)^{-1}\mathbf{A_{2}A_{4}}^{-1}\end{bmatrix}}.}$
Setting $\bm{P}=\bm{\Upsilon}_{j-1},\bm{A}_{1}=\bm{\Gamma}(0),\bm{A}_{2}=\bm{\xi}_{j-1}^{\mathrm{T}},\bm{A}_{3}=\bm{\xi}_{j-1},\bm{A}_{4}=\bm{\Upsilon}_{j-2}$, we have $\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}=\{\bm{C}_{j-1}^{-1}\bm{U}_{j}\bm{V}_{j},\bm{\Upsilon}_{j-2}^{-1}\bm{\kappa}_{j-1}^{\mathrm{T}}-\bm{\Upsilon}_{j-2}^{-1}\bm{\xi}_{j-1}^{\mathrm{T}}\bm{C}_{j-1}^{-1}\bm{U}_{j}\bm{V}_{j}\}$ since $\bm{\Gamma}(j)=\bm{U}_{j}\bm{V}_{j}^{\mathrm{T}}+\bm{\xi}_{j-1}\bm{\Gamma}^{-1}_{j-2}\bm{\kappa}_{j-1}$
Since, $\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{op}},\|\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\leq 1$ and $\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\xi}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\xi}_{2,j}^{\mathrm{T}}\|^{2}_{F}=\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}$

### v2

@@ A.1 Proofs of the Auxiliary Lemmas
We first compute a bound for VAR(1) case. Then use it for VAR(p). From Lütkepohl (2005), we get
$\bm{\Gamma}(j)=\sum_{i=0}^{\infty}\bm{A}^{j+i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}=\bm{A}^{j}\sum_{i=0}^{\infty}\bm{A}^{i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}$
for a VAR(1) process, where $\bm{A}=\bm{\Gamma}(1)\bm{\Omega}$.
Using the telescopic sum setting $\bm{\Gamma}(0)-\bm{A}\bm{\Gamma}(0)\bm{A}=\bm{C}_{1}$, we have
$\sum_{i=0}^{\infty}\bm{A}^{i}\bm{C}_{1}(\bm{A}^{i})^{\mathrm{T}}=\bm{\Gamma}(0)-\lim_{k\rightarrow\infty}\bm{A}^{k}\bm{\Gamma}(0)(\bm{A}^{k})^{\mathrm{T}}=\bm{\Gamma}(0)$
under our parametrization. Hence, $\bm{\Gamma}(j)=\bm{A}^{j}\bm{\Gamma}(0)$.
As $\bm{\Gamma}(1)=\bm{U}_{1}\bm{V}_{1}^{\mathrm{T}}$ and $\bm{C}_{1}=\bm{\Omega}^{-1}-\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}}$, we have $\bm{V}_{1}\bm{\Omega}\bm{V}_{1}^{\mathrm{T}}=\bm{I}$, and hence $\bm{\Omega}^{-1}\geq\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}}$. Alternatively, $\lambda_{\min}(\bm{\Omega}^{-1}-\bm{U}_{1}\bm{U}_{1}^{\mathrm{T}})\geq 0$.
Since, $\bm{V}_{1}^{\mathrm{T}}\bm{\Omega}\bm{V}_{1}=\bm{I}$, we have $\|\bm{V}_{1}^{\mathrm{T}}\bm{\Omega}^{1/2}\|^{2}_{\mathrm{op}}=1$. Hence $\|\bm{V}_{1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Omega}^{-1/2}\|^{2}_{\mathrm{op}}$.
Note that for any two matrices, $\bm{A}^{k}-\bm{B}^{k}=\bm{A}^{k-1}(\bm{A}-\bm{B})+(\bm{A}^{k-1}-\bm{B}^{k-1})\bm{B}$. Recursively using this relation and applying norm, we obtain
$\|\bm{A}^{k}-\bm{B}^{k}\|^{2}_{\mathrm{F}}\leq\|\bm{A}-\bm{B}\|^{2}_{\mathrm{F}}(\sum_{i=1}^{k}\|\bm{A}\|_{\mathrm{op}}^{2(k-i)}\|\bm{B}\|_{\mathrm{op}}^{2(i-1)}),$
using the triangle inequality and the submultiplicative property of operator norm.
After some simplification using GP-series sum results,
$\sum_{i=1}^{k}\|\bm{A}\|_{\mathrm{op}}^{2(k-i)}\|\bm{B}\|_{\mathrm{op}}^{2(i-1)}$ $=\frac{\|\bm{A}\|_{\mathrm{op}}^{2k}-\|\bm{B}\|_{\mathrm{op}}^{2k}}{\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}}$
=$=\frac{\{\|\bm{B}\|^{2}_{\mathrm{op}}+(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})\}^{k}-\|\bm{B}\|_{\mathrm{op}}^{2k}}{\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}}$
$=\sum_{i=0}^{k-1}{k\choose i}\|\bm{B}\|^{i}_{\mathrm{op}}(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})^{k-1-i}.$
Hence,
$\|\bm{A}^{k}-\bm{B}^{k}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{A}-\bm{B}\|^{2}_{\mathrm{F}}\left\{\sum_{i=0}^{k-1}{k\choose i}\|\bm{B}\|^{i}_{\mathrm{op}}(\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}})^{k-1-i}\right\}$
$\leq k\|\bm{A}-\bm{B}\|^{2}_{\mathrm{F}}\min\{\|\bm{A}\|_{\mathrm{op}}^{2(k-1)},|\|\bm{A}\|^{2}_{\mathrm{op}}-\|\bm{B}\|^{2}_{\mathrm{op}}|^{k-1}\}.$
Hence
$\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ $\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}}+\sum_{k=1}^{T-1}(T-k)\|\bm{\Upsilon}_{1}(k)-\bm{\Upsilon}_{2}(k)\|^{2}_{\mathrm{F}};$
$\|\bm{\Upsilon}_{1}(k)-\bm{\Upsilon}_{2}(k)\|^{2}_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}^{k}-\bm{A}_{2}^{k}\|_{\mathrm{F}}+\|\bm{A}_{2}\|^{k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}};$
$\|\bm{A}_{1}-\bm{A}_{2}\|_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1}(1)\|^{2}_{\mathrm{op}}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}+\|\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(1)-\bm{\Upsilon}_{2}(1)\|^{2}_{\mathrm{F}}$
$\leq\|\bm{U}_{1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1}\|^{2}_{\mathrm{op}}\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}$
$\quad+\|\bm{\Omega}_{2}\|^{2}_{\mathrm{op}}\{\|\bm{U}_{1,1}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,1}-\bm{V}_{1,2}\|^{2}_{\mathrm{F}}+\|\bm{V}_{1,2}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,1}-\bm{U}_{1,2}\|^{2}_{\mathrm{F}}\},$
and
$\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$
$\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+\sum_{k=1}^{T-1}(T-k)\{\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}^{k}-\bm{A}_{2}^{k}\|^{2}_{\mathrm{F}}+\|\bm{A}_{2}\|^{k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}\}$
$\leq T\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|_{\mathrm{F}}+\sum_{k=1}^{T-1}(T-k)\{\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}k\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{\mathrm{F}}a^{k-1}$
$\quad+\|\bm{A}_{2}\|^{2k}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}\}$
$\leq(T\frac{b-b^{\mathrm{T}}}{1-b}-\frac{b\{(T-1)b^{\mathrm{T}}-Tb^{T-1}+1\}}{(1-b)^{2}}+T)\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\qquad+T\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{\mathrm{F}}\sum_{k=1}^{T-1}ka^{k-1}$
$\leq(\frac{Tb-Tb^{2}+b^{T+1}-b}{(1-b)^{2}}+T)\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+T\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{A}_{1}-\bm{A}_{2}\|^{2}_{\mathrm{F}}\frac{1+(T-2)a^{T-1}-(T-1)a^{T-2}}{(1-a)^{2}}$ (A.9)
where $a=\|\bm{A}_{1}\|^{2}_{\mathrm{op}}$ and $b=\|\bm{A}_{2}\|^{2}_{\mathrm{op}}$. Since, $a<1$, we have
$\frac{1+(T-2)a^{T-1}-(T-1)a^{T-2}}{(1-a)^{2}}<\frac{1}{(1-a)^{2}}.$
This comes in terms of the model parameters.
@@
It is easy to verify that the functions $(1-a)^{-2}\{1+(T-2)a^{T-1}-(T-1)a^{T-2}\}$ and $(1-b)^{-2}\{Tb-Tb^{2}+b^{T+1}-b\}$ are increasing in $a$ and, $b$ respectively. Following are the limits,
@@
Due to a VAR(1) representation of VAR$(p)$ (Ghosh et al., 2019), Equation (A.9) and $\bm{A}_{\ell}=\bm{\kappa}_{\ell,p}\bm{\Upsilon}_{\ell,p-1}^{-1}$ for $\ell=1,2$, we have
@@
$\leq T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}+T^{3}\|\bm{\Upsilon}_{1,p-1}\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{\mathrm{F}}$
$\leq T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}+T^{3}\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{\mathrm{F}},$
completing the proof. ∎